OpenStudy (anonymous):

Integrate (1/sqrtx . lnx) dx

OpenStudy (raden):

integral by part.. let u = ln(x), du = .... ? dv = 1/sqrt(x) = x^-1/2, v= .... ? so, int (u dv) = uv - int(v du)

OpenStudy (anonymous):

ok,I got it,thx ^^

OpenStudy (raden):

yw :)